# Transverse diffeomorphism invariant spin-2 theories via gauge invariants

### v1

@@ 3.2 NSDiff, NSWTDiff and NSTDiff gauge invariants
Noticing the identity $\partial^{i}\partial^{j}\mathcal{I}_{ij}=0$, it becomes clear that the number of NSDiff invariants $\{\mathcal{I}_{ij},\mathcal{I}_{j},{\cal I}_{00}\}$, i.e., $(D-1)D/2-1+(D-1)+1=D(D+1)/2-1$ matches (7), i.e.,
@@
~$\tilde{\mathcal{I}}_{j}^{t}={\theta_{jk}}(\partial_{i}e_{ik}-\dot{e}_{0k})\,,\qquad I=\theta_{ij}h_{ij}\ ,\qquad I_{ij}^{tt}=\left[\theta_{ik}\theta_{jl}-\frac{\theta_{ij}\theta_{kl}}{(D-2)}\right]h_{kl}\ .$ (43)
@@
At $c=0$ we only have a massless spin-2 particle. Assuming $c\neq 0$, in order to diagonalize the $SO(D-1)$ scalar sector, we can redefine some invariants,
@@
where $(\Psi_{ij}^{tt},\beta_{ij}^{t})=(\Psi_{ji}^{tt},-\beta_{ji}^{t})$ and $\delta_{ij}\Psi_{ij}^{tt}=\partial_{i}\Psi_{ij}^{tt}=\partial_{i}\beta_{ij}^{t}=\partial_{i}\gamma_{i}^{+}=\partial_{i}\gamma_{i}^{-}=\partial_{i}\Phi_{i}^{+}=\partial_{i}\Phi_{i}^{-}=0$. We only need to focus on the $SO(D-1)$ scalars and vectors:
@@
[¯$\begin{bmatrix}\bar{\cal I}_{00}\\ I\\ \bar{I}\\ h\end{bmatrix}=\mathbb{M}_{NTD}\begin{bmatrix}\rho\\ \Lambda\\ \Gamma^{-}\\ \psi\end{bmatrix}-\begin{bmatrix}\dot{\Gamma}^{+}\\ 0\\ 0\\ 0\end{bmatrix}\,,$ (53)
@@
Since $det[\mathbb{M}_{NTD}]=(D-2)\nabla^{4}\,$ the four scalar invariants are independent quantities. Similarly, in the $SO(D-1)$ vector sector, the determinant of the relevant matrix that takes us from $\{\gamma_{j}^{-},\Phi_{j}^{+}\}$ to $\{I_{j}^{t},\tilde{\mathcal{I}_{j}}^{t}\}$ is given by $\nabla^{2}$, so we have again a nonsingular change of variables. In the appendix, we alternatively confirm the particle content of the NSTDiff model (31) via an auxiliary vector field without using Bardeen variables.
Now, when going from the NSDiff to the NSWTDiff basis, we recall that ¯$\bar{\cal I}_{00}$, $\bar{\cal I}$ and $\mathcal{I}$ are the only Weyl noninvariant Bardeen variables. Using the transformation $h_{\mu\nu}\to\bar{h}_{\mu\nu}(h)=h_{\mu\nu}-\eta_{\mu\nu}\,h/D$ to obtain the NSWTDiff invariants, we reach
¯$\bar{\cal I}_{00}^{WT}=\left[\frac{2c-(D-2)}{2cD}\right]e_{00}+\left[\frac{(D-2)+2c(D-1)}{2cD}\right]e_{jj}-\frac{\partial_{j}\dot{e}_{0j}}{\nabla^{2}}-\frac{(1+2c)}{2c}\theta_{ij}e_{ij}\,;$ (55)
¯$\bar{\cal I}_{WT}=\frac{\partial_{j}e_{j0}}{\nabla^{2}}+\left[\frac{(D-2)-2c}{2cD}\right]\frac{\dot{e}_{jj}}{\nabla^{2}}-\left[\frac{2c(D-1)+(D-2)}{2cD}\right]+\frac{\dot{e}_{00}}{\nabla^{2}}-\frac{\theta_{ij}\dot{e}_{ij}}{2c\nabla^{2}}\,;$ (56)
@@
¯$\bar{\cal I}_{00}^{WT}=\left[\frac{2c-(D-2)}{2cD}\right]\rho+\left[\frac{(c+1)(D-2)}{cD}\right]\nabla^{2}\Lambda+\left[\frac{(D-2)+2c(D-1)}{2cD}\right]\nabla^{2}\psi\,;$ (58)
¯$\bar{\cal I}_{WT}=-\left[\frac{2c(D-1)+(D-2)}{2cD}\right]\dot{\rho}-\left[\frac{(D-2)(c+1)}{cD}\right]\dot{\Lambda}+\Gamma^{-}+\left[\frac{(D-2)-2c}{2cD}\right]\dot{\psi}\,;$ (59)
@@
[¯$\begin{bmatrix}\bar{\cal I}_{00}^{WT}\\ \bar{\cal I}_{WT}\\ I_{WT}\\ h\end{bmatrix}=\mathbb{M}_{NSW}\begin{bmatrix}\rho\\ \Lambda\\ \Gamma^{-}\\ \psi\end{bmatrix}-\begin{bmatrix}\dot{\Gamma}^{+}\\ 0\\ 0\\ 0\end{bmatrix},$ (61)
@@ 4 Coupling to external sources
Finally, up to the contact term, the NLDiff model (70) obtained after integration over the antisymmetric components of $e_{\mu\nu}$ may be considered as a linearized version about flat space of the following nonlocal model:
@@
where $R$ is the usual scalar curvature while $C_{j}(-g)\,\,,j=1,2,3$ are functions of the metric determinant $g$ analytic at $g=-1$ such that: $C_{1}(1)>0$, $C_{1}^{\prime}(1)=0$, $C_{2}(1)=c$ and $C_{3}(1)=-B\,C_{1}(1)/4$. This can be checked by performing a weak field expansion according to $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}/\sqrt{C_{1}(1)})$. The action corresponding to (73) is invariant under volume preserving diffeomorphisms. Infinitesimally it amounts to $x^{\mu}\to x^{\mu}+\epsilon^{\mu}$ with $\nabla_{\mu}\epsilon^{\mu}=0$. The metric determinant behaves as a scalar field under such transformations, see comments in [5] and [8]. In the latter reference, a sub case of (73) where the nonlocal term on the curvature is absent, i.e. an ST model with a spin-2 and only one scalar mode, has been investigated phenomenologically. It might be interesting to “localize” the nonlinear SST theory (73) and search for phenomenological effects. This is among our goals.
@@ 5 Conclusion
Here we have investigated the most general Lorentz invariant massless free model of second order in derivatives, described by a rank-2 tensor with both symmetric and antisymmetric parts and invariant under transverse diffeomorphisms ($\partial^{\mu}\epsilon^{T}_{\mu}=0$). By demanding a healthy massless spin-2 particle in the spectrum, we have found three classes of models given in (24), (31) and (32), each parametrized by two real parameters. Except for special points in the parameter’s space where the spectrum is reduced to previously known models in the literature, they all feature two massless scalars (spin-0) and one massless spin-2 (tensor) particle in their spectrum which we call an SST model. There is always a continuous family of SST models such that all three particles are physical.
@@
Integrating over the antisymmetric part of the rank-2 field in the case of the SST models (31), we have obtained a nonlocal model whose natural nonlinear completion is given by the covariant theory (73), where the restrictions on the functions $C_{j}(-g)\,,j=1,2,3$, are given right after (73). Although turning a nonlocal field theory into a local one via auxiliary fields is a nontrivial issue, see for instance [30], one might think of replacing $C_{3}(-g)\sqrt{-g}R(1/\Box)R$ by $C_{3}(-g)\sqrt{-g}(2\,R\,\psi+\psi\,\Box\,\psi)$ in (73) and investigating the phenomenological consequences of a generalization of the gravitational scale invariant theory studied in [8] which corresponds to $C_{3}(-g)=0$.
@@
Finally, the gauge invariant approach adopted here, rooted in Bardeen variables [19], makes the identification of the physical content of the whole model straightforward. Although Lorentz covariance is not manifest, it can be shown that the final theory is Lorentz invariant. We believe that the method may be especially useful in higher rank tensor theories where lower spin (traces) components complicate conventional analyses.
@@ 7 Appendix - NSTDiff content via auxiliary field
In this case we have $q+1<1$. There are three sub cases. If $q+1=0$ the condition $K>0$ leads to $c>0$ then it follows that $f_{D}<0$ which violates the second condition (A-9). If $10<q+1<$ 1, the condition $K>0$ leads to either $c>0$ or $c<-c_{D}/(q+1)$. The case $c>0$ must be ruled out, it leads to $f_{D}<0$ again. If $c<-c_{D}/(q+1)$ we have $c<-c_{D}$ due to $10<q+1<$ 1, it leads once more to $f_{D}<0$ and must be dicarded too. We are left with $q+1<0$. In this case $K>0$ requires $0<c<-c_{D}/(q+1)$, since $f_{D}>0$ demands now $c>-c_{D}/(q+1)$ or $c<-c_{D}$, so there is no solution.

### v2

@@ 3.2 NSDiff, NSWTDiff and NSTDiff gauge invariants
Noticing the identity $\partial^{i}\partial^{j}\mathcal{I}_{ij}=0$, it becomes clear that the number of NSDiff invariants $\{\mathcal{I}_{ij},\mathcal{I}_{j},{\cal I}_{00}\}$, i.e., (D1)D/21+(D1)+1=D(D+1)/21(D-1)D/2-1+(D-1)+1=D(D+1)/2-1 matches (7), i.e.,
@@
$\tilde{\mathcal{I}}_{j}^{t}={\theta_{jk}}(\partial_{i}e_{ik}-\dot{e}_{0k})\,,\qquad I=\theta_{ij}h_{ij}\ ,\qquad I_{ij}^{tt}=\left[\theta_{ik}\theta_{jl}-\frac{\theta_{ij}\theta_{kl}}{(D-2)}\right]h_{kl}\ .$ (43)
@@
At $c=0$ we only have a massless spin-2 particle. Assuming $c\neq 0$, in order to diagonalize the SO(D1)SO(D-1) scalar sector, we can redefine some invariants,
@@
where $(\Psi_{ij}^{tt},\beta_{ij}^{t})=(\Psi_{ji}^{tt},-\beta_{ji}^{t})$ and $\delta_{ij}\Psi_{ij}^{tt}=\partial_{i}\Psi_{ij}^{tt}=\partial_{i}\beta_{ij}^{t}=\partial_{i}\gamma_{i}^{+}=\partial_{i}\gamma_{i}^{-}=\partial_{i}\Phi_{i}^{+}=\partial_{i}\Phi_{i}^{-}=0$. We only need to focus on the SO(D1)SO(D-1) scalars and vectors:
@@
$\begin{bmatrix}\bar{\cal I}_{00}\\ I\\ \bar{I}\\ h\end{bmatrix}=\mathbb{M}_{NTD}\begin{bmatrix}\rho\\ \Lambda\\ \Gamma^{-}\\ \psi\end{bmatrix}-\begin{bmatrix}\dot{\Gamma}^{+}\\ 0\\ 0\\ 0\end{bmatrix}\,$, (53)
@@
Since $det[\mathbb{M}_{NTD}]=(D-2)\nabla^{4}\,$ the four scalar invariants are independent quantities. Similarly, in the SO(D1)SO(D-1) vector sector, the determinant of the relevant matrix that takes us from $\{\gamma_{j}^{-},\Phi_{j}^{+}\}$ to $\{I_{j}^{t},\tilde{\mathcal{I}_{j}}^{t}\}$ is given by $\nabla^{2}$, so we have again a nonsingular change of variables. In the appendix, we alternatively confirm the particle content of the NSTDiff model (31) via an auxiliary vector field without using Bardeen variables.
Now, when going from the NSDiff to the NSWTDiff basis, we recall that $\bar{\cal I}_{00}$, $\bar{\cal I}$ and $\mathcal{I}$ are the only Weyl noninvariant Bardeen variables. Using the transformation $h_{\mu\nu}\to\bar{h}_{\mu\nu}(h)=h_{\mu\nu}-\eta_{\mu\nu}\,h/D$ to obtain the NSWTDiff invariants, we reach
$\bar{\cal I}_{00}^{WT}=\left[\frac{2c-(D-2)}{2cD}\right]e_{00}+\left[\frac{(D-2)+2c(D-1)}{2cD}\right]e_{jj}-\frac{\partial_{j}\dot{e}_{0j}}{\nabla^{2}}-\frac{(1+2c)}{2c}\theta_{ij}e_{ij}\,;$ (55)
$\bar{\cal I}_{WT}=\frac{\partial_{j}e_{j0}}{\nabla^{2}}+\left[\frac{(D-2)-2c}{2cD}\right]\frac{\dot{e}_{jj}}{\nabla^{2}}-\left[\frac{2c(D-1)+(D-2)}{2cD}\right]+\frac{\dot{e}_{00}}{\nabla^{2}}-\frac{\theta_{ij}\dot{e}_{ij}}{2c\nabla^{2}}\,;$ (56)
@@
$\bar{\cal I}_{00}^{WT}=\left[\frac{2c-(D-2)}{2cD}\right]\rho+\left[\frac{(c+1)(D-2)}{cD}\right]\nabla^{2}\Lambda+\left[\frac{(D-2)+2c(D-1)}{2cD}\right]\nabla^{2}\psi\,;$ (58)
$\bar{\cal I}_{WT}=-\left[\frac{2c(D-1)+(D-2)}{2cD}\right]\dot{\rho}-\left[\frac{(D-2)(c+1)}{cD}\right]\dot{\Lambda}+\Gamma^{-}+\left[\frac{(D-2)-2c}{2cD}\right]\dot{\psi}\,;$ (59)
@@
$\begin{bmatrix}\bar{\cal I}_{00}^{WT}\\ \bar{\cal I}_{WT}\\ I_{WT}\\ h\end{bmatrix}=\mathbb{M}_{NSW}\begin{bmatrix}\rho\\ \Lambda\\ \Gamma^{-}\\ \psi\end{bmatrix}-\begin{bmatrix}\dot{\Gamma}^{+}\\ 0\\ 0\\ 0\end{bmatrix}$, (61)
@@ 4 Coupling to external sources
Finally, suppressing the sources, although there is in general no unique non linear completion of linearized theories, the NLDiff model (70) obtained after integration over the antisymmetric components of $e_{\mu\nu}$ may be considered as a linearized version about the flat Minkowiski space of the following nonlocal model:
@@
where $R$ is the usual scalar curvature while $C_{j}(-g)\,\,,j=1,2,3$ are functions of the metric determinant $g$ analytic at g=1g=-1 such that: $C_{1}(1)>0$, $C_{1}^{\prime}(1)=0$, $C_{2}(1)=c$ and $C_{3}(1)=-B\,C_{1}(1)/4$. This can be checked [^10] Recall that (70) at $B=0=c$, without sources, corresponds to twice the linearized Einstein-Hilbert theory $\sqrt{-g}R$ about the flat Minkowiski space. Moreover, $g=-1-h+{\cal O}(h^{2})$ and since the scalar curvature is of first order in $h_{\mu\nu}$, the non local term in (73) becomes the non local term of (71) at quadratic order. by performing a weak field expansion according to $g_{\mu\nu}=\eta_{\mu\nu}+h_{\mu\nu}/\sqrt{C_{1}(1)})$. The action corresponding to (73) is invariant under volume preserving diffeomorphisms. Infinitesimally it amounts to $x^{\mu}\to x^{\mu}+\epsilon^{\mu}$ with $\nabla_{\mu}\epsilon^{\mu}=0$. The metric determinant behaves as a scalar field under such transformations, see comments in [5] and [8]. In the latter reference, a sub case of (73) where the nonlocal term on the curvature is absent, i.e. an ST model with a spin-2 and only one scalar mode, has been investigated phenomenologically. It might be interesting to “localize” the nonlinear SST theory (73) and search for phenomenological effects. This is among our goals.
@@ 5 Conclusion
Here we have investigated the most general Lorentz invariant massless free model of second order in derivatives, described by a rank-2 tensor with both symmetric and antisymmetric parts and invariant under transverse diffeomorphisms ($\partial^{\mu}\epsilon^{T}_{\mu}=0$). By demanding a healthy massless spin-2 particle in the spectrum, we have found three classes of models given in (24), (31) and (32), each parametrized by two real parameters. They are displayed in Table 1. Except for special points in the parameter’s space where the spectrum is reduced to previously known models in the literature, they all feature two massless scalars (spin-0) and one massless spin-2 (tensor) particle in their spectrum which we call an SST model. There is always a continuous family of SST models such that all three particles are physical.
@@
An important direction to follow is to search for non linear completions of the consistent free models of table 1. As a first step, we are now using (work in progress) the Noether iterative method, see, e.g., [30], in order to investigate the construction of consistent interacting cubic vertices along with extensions of the linearized gauge symmetry, similar in spirit to the Gupta program which starts with a spin-2 linearized model and aims to arrive at a full non linear gravitational model via an iterative procedure, see [31] and [32], with the difference that now we are dealing with a rank-2 tensor with an antisymmetric part. On the other hand, one can think of integrating out the antisymmetric part of the tensor as in the case of the SST models (31). So we have obtained a nonlocal model whose natural nonlinear completion is given by the covariant, under volume preserving diffeomorphisms, theory (73) where the restrictions on the functions $C_{j}(-g)\,,j=1,2,3$, are given right after (73). Although turning a nonlocal field theory into a local one via auxiliary fields is a nontrivial issue, see for instance [33], one might think of replacing $C_{3}(-g)\sqrt{-g}R(1/\Box)R$ by $C_{3}(-g)\sqrt{-g}(2\,R\,\psi+\psi\,\Box\,\psi)$ in (73) and investigating the phenomenological consequences of a generalization of the gravitational scale invariant theory studied in [8] which corresponds to $C_{3}(-g)=0$.
@@
We remark that the gauge invariant approach adopted here, rooted in Bardeen variables [19], makes the identification of the physical content of the whole model straightforward. Although Lorentz covariance is not manifest, it can be shown that the final theory is Lorentz invariant. We believe that the method may be especially useful in higher rank tensor theories where lower spin (traces) components complicate conventional analyses. Moreover, recently one has shown [34] the equivalence of different points in the parameters space of some spin-2 and spin-3 models invariant under transverse differomorphisms. We believe that by means of redefinitions of gauge invariants and sources we can achieve a simpler proof of a similar equivalence regarding the parametric families of models in table 1.
Finally, It would be interesting also to figure out how the reduction in the number of effective field components stemming from the use of gauge invariants is connected with the absence of redundant field components in some first order wave equations approach to higher spin particles as in [35].
@@ 7 Appendix - NSTDiff content via auxiliary field
In this case we have $q+1<1$. There are three sub cases. If $q+1=0$ the condition $K>0$ leads to $c>0$ then it follows that $f_{D}<0$ which violates the second condition (A-9). If $0<q+1<1$, the condition $K>0$ leads to either $c>0$ or $c<-c_{D}/(q+1)$. The case $c>0$ must be ruled out, it leads to $f_{D}<0$ again. If $c<-c_{D}/(q+1)$ we have $c<-c_{D}$ due to $0<q+1<1$, it leads once more to $f_{D}<0$ and must be dicarded too. We are left with $q+1<0$. In this case $K>0$ requires $0<c<-c_{D}/(q+1)$, since $f_{D}>0$ demands now $c>-c_{D}/(q+1)$ or $c<-c_{D}$, so there is no solution.